OpenStudy (cutiecomittee123):

Find the remainder of the division problem shown below (x^4+x^3-3x-3)/(x+2)

OpenStudy (freckles):

can you solve x+2=0?

OpenStudy (freckles):

The remainder theorem says if you have \[\frac{P(x)}{x+r} \text{ then the remainder is } P(-r)\]

OpenStudy (cutiecomittee123):

So then x+2=0 results in x=-2

OpenStudy (cutiecomittee123):

does synthetic division work in this case?

OpenStudy (freckles):

yes to that but all you have to do now is evaluate (-2)^4+(-2)^3-3(-2)-3

OpenStudy (cutiecomittee123):